40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Aug 27 13:55:43 2008, rana,jenne, Update, PSL, PMC Servo Board 
    Reply  Fri Aug 29 02:40:43 2008, rana,jenne, Update, PSL, PMC Servo Board pmclp-07.pdfpmclp_40m_080824.mpmclp.m
Message ID: 892     Entry time: Wed Aug 27 13:55:43 2008     Reply to this: 895
Author: rana,jenne 
Type: Update 
Category: PSL 
Subject: PMC Servo Board 
Board is back in. PMC is locked.

Nominal gain is now 15 dB with brick. We need to do more studies:

  • Find out why there is still 35 MHz signal at the error point. Order some low pass filters to cut off above 35 MHz.
  • Explore brick + no-brick loop shapes and error spectra.
  • Measure and set the OLG.

We've left the copper-wrapped lead brick installed to let it slowly conform to the glass better.
ELOG V3.1.3-